A machine has three interlocking gears . Gear A has 18 teeth and turns clockwise at 30 revolutions per minute. Gear B turns counter-clockwise and has 36 teeth. Gear C turns clockwise and has 9 teeth. How many revolutions does gear C turn in one minute ?
Found 3 solutions by Alan3354, stanbon, drk:
Answer by Alan3354(69443)

Gear C has 1/2 as many teeth, so it makes 60 rpm.
The size of the middle gear is not relevant.

Answer by stanbon(75887)

A machine has three interlocking gears .
Gear A has 18 teeth and turns clockwise at 30 revolutions per minute.
Gear B turns counter-clockwise and has 36 teeth.
Gear C turns clockwise and has 9 teeth.
------------------------------------------------------
How many revolutions does gear C turn in one minute ?
In one min a turns 30 revs = 30*18 teeth = 540 teeth
---
In the same min B turns 540/36 = 15 revs = 540 teeth
---
In the same min C turns 540/9 = 60 revs
==========================================
Cheers,
Stan H.

Answer by drk(1908)

Gear A: 18 teeth -> 30 rev / min -> 60 rev / 2 min.
Gear B: 36 teeth
Gear C: 9 teeth.
---
The LCM (18, 36, 9) = 36
9*4 = 36
4*30 = 120
--
So, gear C turns 120 revolutions / min.